# NeuroStrata: Harnessing Neurosymbolic Paradigms for Improved Design, Testability, and Verifiability of Autonomous CPS

### v1

@@ Abstract.
Ivan Ruchkin is supported in part by the NSF Grant CCF-2403616. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation (NSF) or the United States Government.
copyright: rightsretainedccs: Software and its engineering Software verification and validationccs: Computer systems organization Embedded and cyber-physical systemsccs: Computing methodologies Machine learningccs: Theory of computation Formal languages and automata theory
@@ 1. CONTEXT, MOTIVATION, AND AIMS
The integration of machine learning (ML) into CPS has driven innovations in autonomous vehicles (Waymo, 2023; Tesla Autopilot, 2023; Uber ATG, 2023), delivery drones (Amazon Prime Air, 2023; Google Wing, 2023; Zipline, 2023), and robotic surgeries (Davinci, 2023; Mazor, 2023; Mako, 2023). While ML enhances autonomy and intelligence, its uncertain and brittle nature, as witnessed in softmax-based classifications and regression-based control, undermines traditional formal verification, necessitating novel solutions.
@@
Neurosymbolic approaches combine symbolic reasoning with neural learning, addressing challenges in autonomous CPSs by enabling co-training of neural components with symbolic logic through probabilistic logic programming and differentiable reasoning (Cohen et al., 2017; Manhaeve et al., 2021; Li et al., 2023; Lu et al., 2024). Advanced methods, such as program synthesis using Domain-Specific Languages (DSLs), demonstrate promise by supporting both deterministic and probabilistic programs (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021). However, significant bottlenecks persist, including unseen data in real-world deployments (Fredrikson et al., 2023), multi-sensor fusion challenges (Strickland et al., 2018), and the lack of neural component verification. Determinism at the decision level of perception and mission planning is also missing, which can enable the applicability of decades-long formal verification and testing techniques. Ensuring system-level safety and liveness requires systematic co-design of perception, planning, and control — a critical aspect missing in current methods.
This paper summarizes state-of-the-art neurosymbolic paradigms, using autonomous driving as a case study to highlight gaps in adaptability to unseen environments, multi-sensor complexity, systematic validation of neural components, and determinism at the decision level, critical for reliable autonomous CPS.
To address these issues, we put forward NeuroStrata — a neurosymbolic framework for designing and assuring autonomous CPSs. Our vision integrates neurosymbolic distillation and corner-case test generation using LLMs to enable data-driven specification mining, top-down synthesis of symbolic and neurosymbolic components, and runtime bottom-up adaptation via program induction. This approach evolves symbolic programs dynamically for decision-making in perception and planning/control modules. Building on prior work, we aim to transform the testability and verifiability of autonomous CPSs through the neurosymbolic paradigm.
2. Motivating System and the State-of-the-art
@@
Recent advancements in verifying machine learning models and testing learning-enabled CPS have utilized methods like NNV star sets (Tran et al., 2020), Sherlock (Dutta et al., 2019), Reluplex (Katz et al., 2017), and Branch-and-Bound (Bunel et al., 2018). While promising, these methods often fall short in providing comprehensive safety and liveness guarantees, particularly when dealing with the complexity and scale of the real-world, multi-modal CPS as shown in the motivating system above. System-level testing (Deng et al., 2022a; Tian et al., 2022; Li et al., 2020) and robustness analyses (Tian et al., 2018; Deng et al., 2020; Cai and Koutsoukos, 2020) offer progress but still rely on probabilistic methods to assure probabilistic systems (Cleaveland et al., 2022; Badithela et al., 2023), which can weaken the guarantees (Badings et al., 2023).
In parallel, neurosymbolic paradigms aim to combine neural and symbolic approaches — either through differentiable logic programming to produce differentiable yet deterministic outputs (Li et al., 2023; Raghothaman et al., [n. d.]) or through program induction to synthesize deterministic programs by observing limited input-output pairs (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021). However, they face notable limitations. The differentiable logic programming paradigm struggles with hard-coded mappings between neural component outputs and logic program inputs, as well as the complexity of implementing differentiable logic programming capabilities. These constraints limit its ability to support multi-modal sensors and the sophisticated logics required for autonomous systems, which demand more automated and expressive guidance. Similarly, the program induction paradigm relies on Domain-Specific Languages (DSLs) for guidance, which are often hard-coded and lack the flexibility needed for multi-modal sensor integration and advanced reasoning (Parisotto et al., 2016; Ellis et al., 2021; Bjørner et al., 2023). Furthermore, program induction may not be sufficient to replace all pure AI components in such systems.
@@ 3. NeuroStrata: Our Vision for Hierarchical Neurosymbolic Framework for Autonomous Systems
Modules. At design time, Specification Mining, built on neurosymbolic distillation (Singireddy et al., 2023; Blazek et al., 2024; Abir, 2024), extracts formal safety and liveness specifications from training datasets. To cover more diverse safety and liveness violations and out-of-distribution scenarios beyond existing training data, we leverage recent work using large language models to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023), such as front-facing cameras in vehicles, to generate additional real-world crashes and unusual cases from various angles. These specifications are propagated hierarchically across the system. In the perception stack, a high-level Scene Graph encodes semantic relationships and interactions between objects (e.g., “pedestrian crossing road”), represented as differentiable, adaptable programs that can be verified using formal tools like theorem provers. The middle-level Semantic Map encodes spatial and semantic information such as road layouts and drivable areas, ensuring consistency with the scene graph via symbolic rules. The low-level Sensor Fusion and Signal Processing integrates multi-modal sensor data (e.g., LiDAR, cameras, GPS) while enforcing constraints on accuracy and consistency, leveraging neurosymbolic reasoning for fusion and processing. Similarly, the planning and control stack follows a hierarchical structure. The high-level Global/Mission Planner synthesizes deterministic programs to achieve overall system objectives, verified with formal methods such as theorem proving. The middle-level Local Planner generates short-term trajectories that align with global plans while adapting to local changes, guided by symbolic reasoning. The low-level Actuation Control converts trajectories into control commands (e.g., steering angle, throttle) and ensures compliance with constraints using runtime verification techniques.
@@
Adaptation. During runtime, NeuroStrata dynamically adapts its perception and planning modules to real-time data while maintaining formal specification compliance. For perception, sensor data flows upward through the hierarchy, where outputs from the low-level sensor fusion are validated against middle-level semantic map constraints, and updates propagate to the high-level scene graph. It evolves dynamically using differentiable program induction, compacting, and adapting specifications as needed. For planning and control, high-level mission planners adjust strategies based on changing conditions, while differentiable and adaptable control programs refine global plans and compact themselves in response to system data. Middle- and low-level components, such as local planners and actuation control, remain guided by symbolic reasoning to ensure safety and alignment with global objectives. This integration enables simultaneously adaptable and formally validated behavior throughout the system.
@@ 4. EARLY RESULTS AND FUTURE PLAN
We conducted a preliminary case study to investigate a key Research Question (RQ): “can neurosymbolic reasoning complement neural-network training to align with underlying specifications?”. We also outline our future plans, along with the potential challenges and proposed solutions.
4.1. Assessing neurosymbolic reasoning to align neural network training with specifications
In this study, we investigate the capability of differentiable neurosymbolic reasoning to align perceptual neural networks with specifications. We explore the application of a high-level visual perception system trained by aligning its output with specifications. In this application, the goal is to infer spatio-temporal scene graphs (STSG) from videos (e.g., ones taken by ego-centric cameras), where the scene graphs must align with a given spatio-temporal specification. Figure 3 depicts a specification for a traffic scene which is described in natural language but then formalized into a temporal logic formula. Notice that the specification consists of logical symbols like exists ($\exists$), and ($\wedge$), not ($\neg$¬), and finally ($\Diamond$). A neurosymbolic approach to solving this task comprises of a neural model for STSG extraction and a differentiable symbolic component for aligning the predicted STSG with the given specification. Being differentiable, the loss computed from the alignment process can be used to supervise the neural model. We evaluated our work on three datasets: OpenPVSG (Yang et al., 2023), 20BN-Something-Something (Goyal et al., 2017), and MUGEN (Hayes et al., 2022), each with diverse temporal properties. Our approach outperforms current baselines on downstream tasks while offering explainability [^1] Reference to the full report was anonymized for the review process.. This specification alignment provides high confidence in our top-down synthesis approach and in guiding neural training with our specifications.
@@ 4.1. Assessing neurosymbolic reasoning to align neural network training with specifications
4.2. Future Research Plan
To advance NeuroStrata, we propose a six-step future plan with concrete steps to address challenges at each stage.
First, generating diverse training datasets will leverage recent advancements in model-based testing that utilize LLMs to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023). The multi-faceted challenge lies in ensuring the generated datasets are diverse, representative of real-world scenarios, and capable of addressing edge cases. Solutions can be tailored around prior work by accessing diverse sensor datasets and logs, leveraging advanced multi-modal LLMs, and integrating domain-specific constraints with iterative refinement based on industrial partner feedback.
Second, designing suitable DSLs is essential for capturing hierarchical and semantically rich specifications. These DSLs enable experts to encode operational constraints for sensor fusion, signal processing, and physical actuation control. Challenges include ensuring the DSLs are intuitive for domain experts while expressive enough to handle complex requirements. Solutions involve co-designing DSLs with autonomy and robotics specialists, developing language automation, and designing usable visual interfaces. By providing a bridge between formal methods and practical application, these DSLs empower experts to play an active role in system design.
Third, developing a specification mining module based on neurosymbolic distillation will extract formal safety and liveness specifications from training datasets and LLM interactions. Aligning mined specifications with real-world requirements and handling noisy/hallucinated data are key challenges. Hybrid approaches that combine symbolic reasoning with neural embeddings, as well as active learning techniques, can iteratively refine the mined specifications to ensure accuracy and physical grounding.
Fourth, for design-time synthesis and verification, we will enforce multi-level specifications for perception and planning/control modules, leveraging the hierarchical structure defined in our DSL. Modular architectures will enable scalable top-down synthesis of symbolic and neurosymbolic components, while formal verification ensures compliance with specifications. Parallelized processes and adaptive abstraction techniques will address scalability challenges, ensuring robustness across diverse scenarios and high-dimensional inputs.
Fifth, for runtime adaptation and validation, we will develop mechanisms to dynamically refine symbolic programs for real-world changes while ensuring specification compliance. Inspired by program induction approaches like DreamCoder (Ellis et al., 2021), NeuroStrata will iteratively refine symbolic representations using real-time data. Challenges include maintaining computational efficiency and real-time guarantees. To address these, we will optimize runtime validators, integrate lightweight symbolic reasoning for faster adaptation, and implement efficient runtime verification to ensure reliability and compliance with minimal overhead. These advancements will enable NeuroStrata to adapt to dynamic environments and evolving operational conditions.
Finally, for industrial deployment, NeuroStrata will be applied to autonomous driving systems, delivery drones, cargo drones, and passenger aircraft — as facilitated by our partners. Key challenges include seamless integration into existing systems, adherence to stringent safety standards, and building trust among stakeholders. Solutions include close collaborative projects, iterative deployment in increasingly open environments, and the creation of comprehensive documentation and training programs to facilitate adoption.
@@ 5. CONCLUSION
This paper explores the potential and challenges of neurosymbolic paradigms for designing, testing, and verifying autonomous CPS. We propose NeuroStrata, a framework enabling top-down synthesis of symbolic and neurosymbolic components for perception and planning/control, and bottom-up adaptation of symbolic programs for real-time decisions. Early results validate neural alignment with specifications. We outline challenges and solutions for implementing NeuroStrata, aiming to bridge theoretical advancements and practical applications, transforming autonomous CPS testing and verification in real-world scenarios.

### v2

@@ Abstract.
This work is supported in part by the Australian Research Council (FT240100269, DP210102447, LP190100676), the U.S. National Science Foundation (CCF 2403616, NAIAD 2332744), and the National AI Institute for Edge Computing Leveraging Next Generation Wireless Networks (CNS 2112562).
journalyear: 2025copyright: ccconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728500isbn: 979-8-4007-1276-0/2025/06ccs: Software and its engineering Software verification and validationccs: Computer systems organization Embedded and cyber-physical systemsccs: Computing methodologies Machine learningccs: Theory of computation Formal languages and automata theory
@@ 1. CONTEXT, MOTIVATION, AND AIMS
The integration of machine learning (ML) into cyber-physical systems (CPS) has spurred innovations in autonomous vehicles (Waymo, 2023; Tesla Autopilot, 2023; Uber ATG, 2023), delivery drones (Amazon Prime Air, 2023; Google Wing, 2023; Zipline, 2023), and robotic surgery (Davinci, 2023; Mazor, 2023; Mako, 2023). While ML enhances autonomy, its inherent uncertainty and brittleness—evident in softmax-based classifications and regression-based control—limit the applicability of traditional formal verification, calling for new approaches. Neurosymbolic methods address this gap by combining neural learning with symbolic reasoning, enabling co-training via probabilistic logic programming and differentiable reasoning (Cohen et al., 2017; Manhaeve et al., 2021; Li et al., 2023; Lu et al., 2024). Recent advances, including DSL-based program synthesis, support both deterministic and probabilistic behaviors (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021).
Despite progress, key challenges remain: handling unseen deployment data (Fredrikson et al., 2023), multi-sensor fusion (Strickland et al., 2018), and the absence of neural component verification and decision-level determinism. These limitations hinder the application of mature formal methods to autonomous CPS. Ensuring safety and liveness requires systematic co-design of perception, planning, and control—still lacking in current frameworks.
This paper reviews state-of-the-art neurosymbolic paradigms, using autonomous driving to illustrate limitations in adaptability, sensor integration, component validation, and decision determinism—each critical for reliable CPS operation. To address these, we propose NeuroStrata—a neurosymbolic framework for the design and assurance of autonomous CPSs. It integrates neurosymbolic distillation and LLM-guided test generation for data-driven specification mining, top-down synthesis of hybrid components, and runtime bottom-up adaptation via program induction. This enables dynamic evolution of symbolic programs within perception and control modules, advancing the testability and verifiability of autonomous CPSs.
@@
2. Motivating System and State-of-the-art
@@
Recent tools for verifying neural networks, such as NNV star sets (Tran et al., 2020), Sherlock (Dutta et al., 2019), and Reluplex (Katz et al., 2017), focus primarily on verifying local and input-output properties, which are too narrow to capture system-level safety. Some efforts build on these tools for system-level safety (Ivanov et al., 2021; Geng et al., 2024) but are far from verifying real-world systems. While system-level testing (Deng et al., 2022a; Tian et al., 2022; Li et al., 2020) and robustness analysis (Tian et al., 2018; Deng et al., 2020; Cai and Koutsoukos, 2020) offer broader insights, they often rely on probabilistic reasoning (Cleaveland et al., 2022; Badithela et al., 2023), which weakens formal guarantees (Badings et al., 2023).
In parallel, neurosymbolic paradigms aim to unify neural and symbolic methods. One way to do so is via differentiable logic programming for producing deterministic yet trainable outputs (Li et al., 2023; Raghothaman et al., [n. d.]). Another approach is program induction that synthesizes symbolic programs from limited input-output examples (Parisotto et al., 2016; Bjørner et al., 2023; Ellis et al., 2021).
However, neurosymbolic approaches face non-trivial challenges. Differentiable logic programming often requires hard-coded mappings between neural outputs and logical inputs, and its implementation complexity restricts sensor multi-modality and expressive logics needed in autonomous systems. Likewise, program induction relies on fixed domain-specific languages (DSLs) that lack flexibility for sensor fusion or advanced reasoning (Parisotto et al., 2016; Ellis et al., 2021; Bjørner et al., 2023). Our vision aligns with neurosymbolic reinforcement learning (NS-RL), which combines deep learning with symbolic reasoning for control (Acharya et al., 2023); however, NeuroStrata proposes a broader architecture-level redesign. It supports formal specification, verification, and adaptation across perception, planning, and control, with NS-RL serving as a possible component at lower levels. NeuroStrata enables end-to-end reasoning, design-time synthesis, runtime validation, and system-level verification over properties expressed in rich logics.
Neurosymbolic distillation is increasingly adopted to unify neural learning with symbolic interpretability. For example, Automaton Distillation (Singireddy et al., 2023) maps Q-values from deep RL agents into deterministic finite-state automata for policy transfer, while NUDGE (Delfosse et al., 2023) distills neural policies into weighted first-order logic rules using neurally guided abstraction and differentiable forward reasoning. In contrast, NeuroStrata targets broader system-level specifications across perception, planning, and control, and uniquely integrates such distillation with large language models and active learning via L* to extract and refine signal temporal logic specifications, as demonstrated in our prior work (Zheng et al., 2024).
@@ 3. NeuroStrata: Our Vision for Hierarchical Neurosymbolic Framework for Autonomous Systems
Modules. At design time, Specification Mining, built on neurosymbolic distillation (Singireddy et al., 2023; Blazek et al., 2024; Delfosse et al., 2023), extracts formal safety and liveness specifications from training datasets. To cover more diverse safety and liveness violations and out-of-distribution scenarios beyond existing training data, we leverage recent work using large language models to analyze multi-modal sensor data (Zheng et al., 2024; Deng et al., 2023), such as front-facing cameras in vehicles, to generate additional real-world crashes and unusual cases from various angles. These specifications are propagated hierarchically across the system. In the perception stack, a high-level Scene Graph encodes semantic relationships and interactions between objects (e.g., “pedestrian crossing road”), represented as differentiable, adaptable programs that can be verified using formal tools like theorem provers. The middle-level Semantic Map encodes spatial and semantic information such as road layouts and drivable areas, ensuring consistency with the scene graph via symbolic rules. The low-level Sensor Fusion and Signal Processing integrates multi-modal sensor data (e.g., LiDAR, cameras, GPS) while enforcing constraints on accuracy and consistency, leveraging neurosymbolic reasoning for fusion and processing. Similarly, the planning and control stack follows a hierarchical structure. The high-level Global/Mission Planner synthesizes deterministic programs to achieve overall system objectives, verified with formal methods such as theorem proving. The middle-level Local Planner generates short-term trajectories that align with global plans while adapting to local changes, guided by symbolic reasoning. The low-level Actuation Control converts trajectories into control commands (e.g., steering angle, throttle) and ensures compliance with constraints using runtime verification techniques.
@@
Adaptation. At runtime, symbolic reasoning and neural components interact through a bi-directional feedback loop. Neural modules (e.g., sensor fusion and perception) generate scene-level representations, which are continuously validated against symbolic specifications such as semantic constraints and temporal logic. These neural outputs trigger updates to symbolic programs (e.g., the scene graph program and global planning program) via differentiable program induction (see Figure 2). This progressively evolves symbolic components in response to new observations. In turn, symbolic planners guide neural policy adaptation and ensure specification compliance at all layers. This closed neuro-symbolic loop supports runtime behavior that is both adaptive and formally grounded.
@@ 4. EARLY RESULTS AND FUTURE PLAN
We conducted a preliminary case study to explore the key Research Question (RQ): “Can neurosymbolic reasoning complement neural network training to better align with specifications?” We also outline future directions, challenges, and proposed solutions.
4.1. Specification-Aligned Training
[…]
4.2. Advancing NeuroStrata: A Six-Step Plan
To operationalize NeuroStrata, we propose a six-step roadmap that addresses key challenges in specification-driven learning and adaptation for AI-based CPS.
I. Diverse training dataset generation will leverage model-based scenario synthesis using large language models, as demonstrated in (Zheng et al., 2024; Elmaaroufi et al., 2024), and validated on real systems in (Deng et al., 2023). This approach mitigates covariate shift by generating multi-modal, edge-case–rich training and testing datasets, grounded in physical realism and enhanced with domain-specific constraints.
II. Domain-specific languages (DSLs) will be co-designed with robotics and autonomy experts to express hierarchical specifications across sensing, planning, and control. These DSLs serve as a bridge between formal methods and system engineering, ensuring usability while enabling precise constraint encoding for downstream reasoning and synthesis.
III. Specification mining via neurosymbolic distillation will extract logical properties from training data and LLM queries. To manage hallucinated or noisy outputs, we will integrate symbolic-neural hybrid techniques and active learning to iteratively refine mined specifications for correctness and interpretability.
IV. During design-time synthesis and verification, symbolic and neurosymbolic modules will be synthesized from DSL-defined specifications and verified using compositional and abstraction-based methods. Parallel verification and modular analysis will allow scalability to high-dimensional state spaces, ensuring robust deployment in complex CPS environments.
V. For runtime adaptation and validation, NeuroStrata will employ differentiable program induction (Ellis et al., 2021) to update symbolic programs on-the-fly. To ensure computational feasibility, we will build on the efficient runtime monitoring infrastructure in (Zheng et al., 2016), which uses low-overhead logging and an aspect-oriented programming paradigm to implement scalable runtime verifiers. This setup supports dynamic refinement of neural-symbolic components with minimal latency, addressing both real-time constraints and covariate shift at runtime.
VI. For industrial deployment, NeuroStrata will transition from simulation-based testing in CARLA, and AirSim-based environments to real-world integration with autonomous vehicles, uncrewed aerial vehicles, and robotics platforms. We have initiated research agreements with industry partners in these domains to support this transition. Our upcoming Shonan Seminar (No. 235) and Dagstuhl Seminar (No. 202501048) will further accelerate industry engagement and foster collaborative development and deployment of neurosymbolic CPS solutions.
@@ 5. CONCLUSION
This paper investigates the potential of neurosymbolic paradigms in the design, testing, and verification of autonomous CPS. We introduce NeuroStrata, a framework that supports top-down synthesis of symbolic and neurosymbolic components for perception and control, and bottom-up adaptation of symbolic programs for real-time decision-making. Preliminary results demonstrate neural alignment with specifications. We discuss key challenges and implementation strategies, aiming to bridge theory and practice in autonomous CPS verification.